OpenStudy (anonymous):

Calc question: d^2y/dx^2 of x^2+y^2=25 at point (4,3) ?

OpenStudy (anonymous):

Implicit differentiation to get \(\dfrac{dy}{dx}\), then implicit differentiation to find \(\dfrac{d^2y}{dx^2}\). When you're done with that, plug in \(x=4,y=3\). Do you need help with the differentiating?

OpenStudy (anonymous):

what does implicit mean?

OpenStudy (anonymous):

double derivation ?

OpenStudy (anonymous):

yes i do need help

OpenStudy (anonymous):

Suppose you're given the function \(y=x\). Finding \(\dfrac{dy}{dx}\) is fairly simple, just differentiate both sides with respect to \(x\): \[\frac{d}{dx}[y]=\frac{d}{dx}[x]\\ \frac{dy}{dx}=1\] But now consider the *relation* \(y^2=x\). The difference is that \(y\) is no longer a function of \(x\), but you're still capable of finding the rate of change of \(y\) with respect to \(x\). Implicit differentiation comes into play in this situation; it's essentially using the chain rule: \[\frac{d}{dx}[y^2]=\frac{d}{dx}[x]\\ 2y\frac{d}{dx}[y]=1\\ 2y\frac{dy}{dx}=1\\ \frac{dy}{dx}=\frac{1}{2y}\] If the examples don't make it clear, you can definitely learn a lot from this site: http://tutorial.math.lamar.edu/Classes/CalcI/ImplicitDiff.aspx

OpenStudy (anonymous):

hmm, i don't see how actually

OpenStudy (anonymous):

Let's try with your question, in that case: \[x^2+y^2=25\] Differentiating both sides (for the first time) gives \[\frac{d}{dx}[x^2+y^2]=\frac{d}{dx}[25]\\ \frac{d}{dx}[x^2]+\frac{d}{dx}[y^2]=0\\ 2x+2y\frac{dy}{dx}=0\\ \frac{dy}{dx}=-\frac{x}{y}\] At the point \((4,3)\), you have \[\frac{dy}{dx}=-\frac{4}{3}\] Did that make sense?

OpenStudy (anonymous):

it is d^2y/dx^2 i don't get